OpenStudy (anonymous):

3x=6+9+2x solve for x

OpenStudy (anonymous):

3x-2x=9+6 x=15

OpenStudy (anonymous):

subtract 2x from the right and the left making it x=6+9, combine the 6 and 9 making it 15, therefore x=15